USING MATHEMATICAL INDUCTION P(n)=1+22+2n+........+2n= 2n+1 - 1

1
Expert's answer
2021-01-12T04:19:08-0500

* I think the question is wrong. The statement will be such as "P(n): 1+2^1+2^2+.......+2^n=2^{n+1}-1"


Let "P(n)" be a statement such that "P(n): 1+2^1+2^2+.......+2^n=2^{n+1}-1"

Then "P(n)" is true for "n=0" as "P(0):1=2^{0+1}-1"

Now we show that the statement is true for smallest natural number "n=0" .

Next we have to show for any "k\geq 0" , if "P(k)" holds, then "P(k+1)" also holds.

Assume that for "n=k" , "P(k)" is true. "i.e" "P(k): 1+2^1+2^2+.......+2^k=2^{k+1}-1" .......(1)

Now we have to prove that "P(k+1)" is true.

Then "1+2^1+2^2+.......+2^k+2^{k+1}=(2^{k+1}-1)+2^{k+1}" [using (1)]

"=2.2^{k+1}-1"

"=2^{(k+1)+1}-1"

Therefore the statement "P(k+1)" holds true.

Hence by the principle of mathematical induction the given statement "P(n)" is true for all "n\in N" .